OpenStudy (anonymous):

Help me with this problem plz.. Q: PCl5 <-----> PCl3 + Cl2 In this reversible reaction, the number of moles of Cl2 produced will be ______? If one mole of PCl5 is heated at 250 degree C in a vessel having capacity of of 10 dm3 (At 250 degree C, Kc= 0.041)

OpenStudy (anonymous):

use the formula for Kc

OpenStudy (anonymous):

i have tried but it didn't give the answer..

OpenStudy (anonymous):

Answer is : 0.465 moles

OpenStudy (anonymous):

show me your calculation

OpenStudy (anonymous):

hmm... Kc=x2/V(a-x) o.o41= x2 /10(1-x) x=???